OpenStudy (anonymous):

AP maths a particle moves along the positive x-axis, at time t seconds after leaving point O from rest, the displacement of the particle from O is x cm. The acceleration is defined by a= {5-3t , 0 < or equal t < or equal 1 -(4x+1), t>1 find (a) the max velocity of the particle and its distance from O when this happens, (b)the speed of the particle when x=5/2.

OpenStudy (anonymous):

a=v(dv/dx) or dv/dt 0r d^(2)x/dt^(2)

OpenStudy (anonymous):

use a=dv/dt for first second and a=v(dv/dx) from later on to find expressions foe velocities now find the maximums in each case. the maximum of the wo gives the maximu velocity

OpenStudy (anonymous):

You should get 1/2 ms^-1 , I think. At x = 1/4m

OpenStudy (anonymous):

ya mistake so i delet the post..

OpenStudy (anonymous):

Yeah. Got it. Didnt refresh the page :P

OpenStudy (anonymous):

for d second equation right??

OpenStudy (anonymous):

a=0 -(4x+1)=0 4x+1=0 x=-1/4

OpenStudy (anonymous):

\[a=5-3t\] \[v=5t-\frac{3}{2}t^2+c\] i guess can we find c?

OpenStudy (anonymous):

c=0 bcoz when x=0, t=0, v=0

OpenStudy (anonymous):

oh starts at rest so i suppose c = 0

OpenStudy (anonymous):

and for \(t> 1\) we have \(a=-(4t+1)=-4t-1\) which confuses me a bit because it seems we have a jump discontinuity in acceleration, a phenomenon rare in nature (nature being by in large continuous at the observable level)

OpenStudy (anonymous):

ya.. what u say is correct.. bcoz my sch teacher always comment that the local book are more on theory only instead of giving real situation

OpenStudy (anonymous):

ok no matter we have \[v(t)=-2t^2-t+c\] and as for c maybe we should at least make sure that the velocity in continuous

OpenStudy (anonymous):

for the second equation.. t>1 when a=-(4x-1) not t.... thn wwe need to use a=v(dv/dt) to solve the equation...

OpenStudy (anonymous):

we*

OpenStudy (anonymous):

wait acceleration is a function of x?

OpenStudy (anonymous):

ya for t>1

OpenStudy (anonymous):

but for t is in between [0,1] we use acceleration as a function of t

OpenStudy (anonymous):

then you need so solve a second order differential equation what class is this?

OpenStudy (anonymous):

you have \(x''(t)=-4x(t)+1\)

OpenStudy (anonymous):

differential equation the application part

OpenStudy (anonymous):

ooooooooooooooooooooooooooooh then i guess we have to solve that one yikes i am sure it is easy and i am sure i don't know how to do it right off the bat any ideas?

OpenStudy (anonymous):

I dont think we have to integrate twice use a=vdv/dx

OpenStudy (anonymous):

ok go nuts. i think it will be some combination of sine and cosine, but since you have the velocity, which is the second derivative, as a function of position, i think you have to sole a second order equation i could easily be wrong

OpenStudy (anonymous):

\[v ^{2}=2 \int\limits_{i}^{f}a dx\]

OpenStudy (anonymous):

i think i can solve the part (a) but it is very long.....

OpenStudy (anonymous):

max velocity is dv/dt=0 ==> a=0 for t is in between [0,1] 0=5-3t t=5/3 but max t=1  let when t=1, v=v1 when t=0,v=v0 therefore v1>v0 therefore for the equation to valid, v1=max v a=5-3t dv/dt=5-3t integrate dv=intergrate 5-3t dt v=5t-3t^(2)/2+c when t=0,x=0,v=0 so c=0 v=5t-3t^(2)/2 at t=1 v = v max = 5(1)-3(1)^(2)/2 =7/2